OpenStudy (anonymous):

Alex set the VCR to record for 4 hours. If the movie ran for 2 hours 47 minutes and there were an additional 67 minutes of commercials, how many extra minutes of tape did the VCR record, if any?

OpenStudy (mathmale):

Recall that 60 minutes = 1 hour. The recorder records the sum of 2 hours 47 minutes and 67 minutes, total. Can you calculate that? Add: 2 hours 47 minutes 67 minutes ------------------ ? Simplify this, remembering (again!) that 60 minutes = 1 hour.

OpenStudy (texaschic101):

I would turn everything to minutes... 1 hr = 60 min 4 hrs = (4 * 60) = 240 minutes 2 hrs and 47 minutes = 120 + 47 = 167 minutes and then there is the other 67 minutes so your problem is : 240 - (167 + 67) = your answer you can do the math